OpenStudy (anonymous):

Which property match with the equation? Properties: Addition property of equality, Subtraction property of equality, Multiplication property of equality, Division property of equality, Simplifying, and Distributive property. Equations: 9u-2 = -81 (subtraction property I think?) 9u-27+27 = -81+27 9u = -54 9u/9 = -54/9 (division property I think?) u = -6

OpenStudy (phi):

the first equation looks like it has a typo

OpenStudy (anonymous):

Ah yes, It's supposed to be 9u-27 = -81 !

OpenStudy (phi):

going from 9u-27+27 = -81 + 27 to the next line 9u= -54 looks like simplifying

OpenStudy (anonymous):

Both look like simplifying?

OpenStudy (phi):

if you started with a line 9(u-3)= -81 then going to 9u-27= -81 would be distributing

OpenStudy (phi):

The question is a bit goofy, because it the the "step" you take to get from one line to the next that uses the property. I guess we label the "result line" with the property. like this 9u+27-27 = -81+27 9u = -54 Simplify

OpenStudy (phi):

can you post your labels for each line?

OpenStudy (anonymous):

Yes. Basically the question is asking which equation matches with which property.

OpenStudy (phi):

I meant, what is your answer? It looks like they start with an equation, and they "solve for u" and they want the reason that allows you to write the next line as you work out the solution

OpenStudy (phi):

For example if you started with a line 9(u-3)= -81 then the reason you are allowed to write the next line as 9u-27= -81 would be distributive property

OpenStudy (anonymous):

I'm not sure how to answer the question then.

OpenStudy (phi):

distribute in English means "hand out" and in math 9(u-3) we can "hand out" the 9 to each term inside the parens (we multiply 9 times each term inside the parens) in other words, 9(u-3) = 9u -3*9 = 9u-27 because we use the distributive property

OpenStudy (phi):

9(u-3)= -81 9u-27 = -81 <-- Distributive Property 9u-27+27 = -81+27 9u = -54 9u/9 = -54/9 (division property I think?) YES u = -6

OpenStudy (phi):

now to get from the second line to the third line: 9u-27 = -81 <-- Distributive Property 9u-27+27 = -81+27 <--- what did we do here ?

OpenStudy (anonymous):

Subtraction property? And the third to fourth is simplify?

OpenStudy (phi):

what is different between this line: 9u-27 = -81 and this line: 9u-27 +27 = -81 +27 ?

OpenStudy (anonymous):

They add 27 to both sides

OpenStudy (phi):

adding 27 to both sides is ok, it keeps the equation true. (if we add 27 to only 1 side, that does not work) To remember that it is ok to add the same thing to both sides, people gave a name to the rule that says it's ok: addition property

OpenStudy (phi):

9(u-3)= -81 9u-27 = -81 <-- Distributive Property 9u-27+27 = -81+27 <-- Addition Property of Equality 9u = -54 <---Simplify 9u/9 = -54/9 (division property I think?) YES u = -6 <-- Simplify

OpenStudy (anonymous):

But we're using simplify twice.

OpenStudy (phi):

Yes. I think this is a problem where you have to show how to get from one line to the next, and you have to pick out the correct reason. In this case, we don't use every reason in the list.

OpenStudy (phi):

and we use the same reason two different times. That's ok